You have a buffer solution composed of 4.00 mol of acid and 8.75 mol of the conjugate base. If the pKa of the acid is 5.00, what is the pH of the buffer?
Answer #1

Given:
pKa = 5

use:
pH = pKa + log {[conjugate base]/[acid]}
which becomes:
pH = pKa + log {mol(conjugate base)/mol(acid)}
= 5+ log {8.75/4}
= 5.34

Answer: 5.34